A machine with a cost of $69,500 has an estimated residual value of $3,853 and an estimated life of 5 years or 17,789 hours. What is the amount of depreciation for the second full year, using the double-declining-balance method?

a. $13,900 b. $27,800 c. $26,259 d. $16,680
Answer #1

Annual depreciation rate as per straight line method=100%/5=20%/year

Hence depreciation as per double decline balance=2*Annual depreciation rate as per straight line method*Beginning value of each period

Year Beginning value Depreciation Ending value
1 69500 (2*20%*69500)=27800 (69500-27800)=41700
2 41700 (2*20%*41700)=$16680

Hence depreciation for the second full year using double-declining-balance method

=$16680